OpenStudy (anonymous):

Is the following operation possible? Why or why not? [2] [5 6 0] [-1]

OpenStudy (anonymous):

Matrix operation?

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

\[\left(\begin{array}{ccc} 5&6&0 \end{array}\right)\left(\begin{array}{ccc} 2 \\ -1 \end{array}\right)\]

OpenStudy (unklerhaukus):

a 1x3 matrix times a 2x1 matrix ?

OpenStudy (anonymous):

By multiplying [2] [5 6 0 ] you will get a 1x3 matrix. Which cannot be multiplied with a 1x1 matrix. You require a 3xn Matrix to multiply with a 1x3 matrix. Hope it helps.

OpenStudy (anonymous):

so it is not possible?

OpenStudy (anonymous):

What do you think?